Paul is 8, exactly 1/4 as old as his father. In how many years will Paul be half as old as his father?
Answer by Jk22(389)

The age of the father is 4*8=32y.

In n years, Paul will be 4+n y. old and the father : 32+n, Paul's age will be half of father's age : 4+n=16+n/2

n/2 = 12, n=24years, this will happen in 24years, when Paul will be 32 years old and the father 64y.old.